A taut string of length 2l , fastened at both ends, is disturbed from its position of equilibrium by imparting to each of its points on an initial velocity of magnitude k(2lx-x^2). Find the displacement function y(x,t)

1
Expert's answer
2022-01-12T08:00:29-0500

The displacement "y(x,t)" of any point '"x" ' of the string at any time '"t" ' is given by

"\frac{\partial²y}{\partial t²}=a²\frac{\partial²y}{\partial x^{2}}" "---(1)"

We have to solve equation (1) satisfying the following boundary conditions.

"y(0,t)=0," "for" "t\u22650" "----(2)"

"y(2l,t)=0," "for" "t\u22650" "----(3)"

"y(x,0)=0," "for" "0\u2264x\u22642l" "----(4)"

"\frac{\delta y}{\delta t}(x,0)=k(2lx-x²)," "for" "0\u2264x\u22642l" "----(5)"

The suitable solution of Eq (1), consistent with the vibration of the string,is

"y(x,t)=(A\cos px + B\sin px)"

"(C\cos pat+D\sin pat)" "----(6)"

Using boundary conditions (2) in (6), we have

"A(C\cos pat+D\sin pat)=0" "for" "all" "t\u22650"

"A=0"

Using boundary conditions (3) in (6), we have

"B\sin 2lp(C\cos pat+D\sin pat)=0"

"for" "all" "t\u22650"

Either "B=0" or "\sin 2lp=0"


If we assume that "B=0," we get a trivial solution.

"\sin 2lp=0"

"2lp=n\u03c0"

"p=\frac{nπ}{2l}"

Where "n=0,1,2,...\infty"

Using boundary conditions (4) in (6), we have

"B\sin px.C=0"

"for" "0\u2264x\u22642l"

As "B\u22600," we get "C=0"

Using these values of "A," "p," "C" in (6), the solution reduces to

"y(x,t)=k\sin \frac{nπ}{2l}\sin \frac{nπat}{2l}"

"----(7)"

where "n=0,1,2,3...\infty"

The most general solution of Eq.(1) is

"y(x,t)=\sum_{n=1}^{\infty}\lambda_{n}\sin \frac{nπ}{2l}\cos \frac{nπat}{2l}" "----(8)"

Differentiating both sides of (8) partially with respect to t, we have

"\frac{\delta y}{\delta t} (x,t)=" "\sum_{n=1}^{\infty}(\frac{nπa}{2l}.\lambda_{n})\sin \frac{nπx}{2l}\cos \frac{nπat}{2l}" "----(9)"

Using boundary condition (5) in (9), we have

"\sum_{n=1}^{\infty}(\frac{nπa}{2l}\lambda_{n})\sin \frac{nπx}{2l}=k(2lx-x²)"

"for" "0\u2264x\u22642l"

"=\sum_{n=1}^{\infty}b_{n}\sin \frac{nπx}{2l}"

Which is Fourier half-range sine series of "k(2lx-x\u00b2)" in "(0,2l)."

Comparing like terms,we get

"\frac{nπa}{2l}.\lambda_{n}=b_n="

"\frac{2}{2l}\int_{0}^{2l}f(x)\sin \frac{nπx}{l}dx," by Euler's formula

"=\frac{2}{2l}\int_{0}^{2l}k(2lx-x²)\sin \frac{nπx}{2l}dx"

"=\frac{2k}{n\pi a}[\,(2lx-x^{2})[\,\frac{-cos\frac{n\pi x}{2l}}{\frac{n\pi}{2l}}]\,-(2l-2x[\,\frac{-sin\frac{n\pi x}{2l}}{\frac{n^{2}\pi^{2}}{4l^{2}}}]\,\\+(-2)[\,\frac{cos\frac{n\pi x}{2l}}{\frac{n^{3}\pi^{3}}{8l^{3}}}]\,]\,_{0}^{2l}"

"=\frac{32kl^{3}}{n^{4}\pi^{4}a}[1-(-1)^{n}]"


"=\begin{cases}0,\hspace{0.4cm}if\hspace{0.1cm} n \hspace{0.1cm}is\hspace{0.1cm} even\\\\\\ \frac{32kl^{3}}{n^{4}\pi^{4} a},\hspace{0.4cm}if\hspace{0.1cm} n \hspace{0.1cm}is\hspace{0.1cm} odd\end{cases}"


Using this value of "\lambda_n" in (8), the required solution is

"y(x,t)=" "\frac{64kl³}{π⁴a}\sum_{n=1}^{\infty}\frac{1}{(2n-1)⁴}\sin \frac{(2n-1)πx}{2l}\cos \frac{(2n-1)πat}{2l}"